3x - 4y = 4
x + 2y = 8
(Solve using any method, and please name the method as well)

Answer by stanbon(75887)

3x - 4y = 4
x + 2y = 8
-----
Method: elimination
Multiply thru the 2nd equation by 2:
3x - 4y = 4
2x + 4y = 16
----
Add and solve for "x":
5x = 20
x = 4
---
Solve for "y":
x + 2y = 8
4 + 2y = 8
2y = 4
y = 2
----
Solution: (4,2)
====================
Cheers,
Stan H.
===========